# DIO: Refining Mutual Information and Causal Chain to Enhance Machine Abstract Reasoning Ability

### v4

@@ V-A The learning objective of the DIO
Correspondingly, the optimization objective of DIO, the variational lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, is updated as follows:
$I\geq\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{\beta-9}\sum_{{\hat{\alpha}}=9,{\hat{\alpha}}\neq\alpha}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right]$ (16)
Intuitively, the looseness of the updated lower bound has been reduced.
However, as shown in Formula (16), if these introduced hypothetical options fail to influence DIO’s decisions, this new lower bound, when used as an optimization objective, will not bring DIO any significant improvement. Specifically, when the probabilities that DIO assigns to the hypothetical options $\{x_{i}|i\in[17,\beta]\}$ approach zero, whereby these options cease to perturb the original distribution $\{\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\alpha}\in[9,16]\}$, the variational lower bounds derived from Equation (16) and Equation (12) become equivalent when employed as loss functions. This demonstrates that unconstructive options, which are readily assigned low weights and low probabilities by DIO, contribute nothing to tightening the lower bound.
Therefore, this paper proposes introducing extra constructive hypothetical options into DIO, options which can reinforce acceptance of the correct options and simultaneously strengthen repulsion of the incorrect ones, so as to steer its learning objective toward a tighter variational lower bound and, in turn, ultimately further enhance its performance.
VI Brando: Tightening the Variational Lower Bound of Mutual Information by Constructing Hypothetical Options.
This paper has put forward the idea that incorporating extra constructive hypothetical options into DIO’s training process offers a promising avenue for boosting its performance. To this end, this section introduces a training framework specifically designed to furnish DIO with such hypothetical options, thereby enhancing its overall effectiveness.
@@
The RPM problem expects the training process of the solving model to involve only a limited number of supervised signals, such as the index of correct options. Relying on manual design to create constructive options goes against this original intent of RPM problems. Therefore, this paper proposes training a set of constructive options instead. This methodology is referred to as Brando (Bounded Random Options Mapping Method).
This paper establishes a set of trainable hypothetical options $\{x_{i}|i\in[17,\beta]\}$ for the context matrix $\{x_{i}|i\in[1,8]\}$. This paper is not concerned with the specific attributes of these options; rather, it focuses solely on how much these options can alter DIO’s evaluation of the original options $\{x_{i}|i\in[9,16]\}$. In other words, we are solely concerned with the constructiveness of these hypothetical options with respect to the context matrix $\{x_{i}|i\in[1,8]\}$ in DIO, rather than their specific configurations. To this end, this paper designs a new loss function $\ell_{\text{Brando}}$ to train these hypothetical options for their constructiveness:
$\ell_{\text{Brando}}=\ell_{\text{up}}+\ell_{down}$ (17)
Where $\ell_{\text{up}}$:
$\ell_{\text{up}}=-\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9,{\hat{\alpha}}\neq\alpha}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (18)
and $\ell_{\text{down}}$:
$\ell_{\text{down}}=-\frac{1}{\beta-17}\sum_{{\hat{\beta}}=17}^{\beta}\log\frac{e^{\text{DIO}(\{x_{i},x_{\hat{\beta}}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9,{\hat{\alpha}}\neq\alpha}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (19)
@@
    It can be observed that $\ell_{\text{up}}$ requires the probability $\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ derived by DIO to be greater than all elements in the set $\{\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|{\hat{\alpha}}\in[9,\beta],{\hat{\alpha}}\neq\alpha\}$. In other words, $\ell_{\text{up}}$ demands that DIO be able to identify the correct option amidst the interference from the hypothetical options $\{x_{i}|i\in[17,\beta]\}$.
@@
    On the other hand, $\ell_{\text{down}}$ constrains that any element in the set $\{\hat{P}(x_{\hat{\beta}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\beta}\in[17,\beta]\}$ must be higher than all elements in the set $\{\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\alpha}\in[9,16],\hat{\alpha}\neq\alpha\}$. In other words, $\ell_{\text{down}}$ ensures hypothetical options $\{x_{i}\mid i\in[17,\beta]\}$ outscore the original incorrect options $\{x_{i}\mid i\in[9,16],i\neq\alpha\}$ under DIO.
In the loss function $\ell_{\text{Brando}}$, the term $\ell_{\text{up}}$ serves the purpose of tightening the variational lower bound, while $\ell_{\text{down}}$ acts as a safeguard to prevent hypothetical options from degenerating into unconstructive ones. By adopting $\ell_{\text{Brando}}$ as the training loss function, DIO not only obtains constructive hypothetical options but also parameterizes a tighter variational lower bound, thereby maximizing the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ more effectively.
@@ VI-B Map Out Hypothetical Options
Evidently, the mapping $\mathcal{F}(\cdot)$ is required to derive a variable number ($\beta-17$) and diverse hypothetical options for the context matrix $\{x_{i}|i\in[1,8]\}$. This indicates that conventional deep neural networks struggle to undertake the task of the mapping $\mathcal{F}(\cdot)$. Firstly, constructing a network with variable and dynamic output specifications is highly challenging. Additionally, from the perspective of neural network optimization, achieving a many-to-many mapping is infeasible. Given these challenges, we need to explore an alternative approach. Therefore, this paper proposes a many-to-one mapping approach:
$x_{\hat{i}}=\hat{\mathcal{F}}(\{x_{i},o_{\hat{i}}|i\in[1,8]\})$ (21)
$\text{Where}\quad o_{\hat{i}}\sim N(\mu,\sigma^{2})$
@@
$z_{\hat{i}j}=\text{Brando}(\{z_{ij},o_{\hat{i}}|i\in[1,8]\})$ (23)
$\text{Where}\quad o_{\hat{i}}\sim N(\mu,\sigma^{2})$
Formula (23) shows that we move away from the option-mapping paradigm and instead adopt a representation-mapping approach, easing the learning pressure on Brando. According to Formula (23), we have designed the structure of the Brando network as shown in the Figure 5.
@@
As indicated in Figure 5, the Brando network is formed by stacking multiple logical layers, with no parameter sharing among each layer. The input to each logical layer consists of $\{z_{ij}|i\in[1,8]\}$ and $o_{\hat{i}}$. Except for the first layer, the input to all subsequent layers comes from the output of the preceding layer. For the first layer, $\{z_{ij}|i\in[1,8]\}$ are the features extracted by DIO from the context matrix, while the $o_{\hat{i}}$ is obtained through the reparameterization trick [2] from a learnable Gaussian distribution $N(\mu,\sigma^{2})$ parameterized by mean $\mu$ and variance $\sigma^{2}$. For the final layer, this paper treats its output $o_{\hat{i}}$ as the representation $z_{\hat{i}j}$ of a hypothetical option $x_{\hat{i}}$ provided by Brando with respect to the context matrix $\{x_{i}|i\in[1,8]\}$.
@@
    After these $Q$, $K$, and $V$ representations are appended positional encoding and processed by layer normalization, the multi-head self-attention mechanism computes interactions among them, yielding the attention outputs $\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$:
@@
    As indicated by the computation step 1 in the Figure 5, the normalized $Q$ representations employ a linear attention mechanism to learn the historical input feature information stored in the memory bank $M_{n-1}$ and buffer $B_{n-1}$, resulting in the memory outputs $\{\tilde{z}_{ij},\tilde{o}_{\hat{i}}|i\in[1,8]\}$. The learning process can be interpreted as:
@@
    Where basis function $\phi(\cdot)$ is an element-wise ELU+1 operation.
@@
    As indicated by the computation steps 2 and 3 in the Figure 5, the memory bank $M_{n-1}$ and buffer $B_{n-1}$ are updated by leveraging the $K$ and $V$ representations derived from the current input features. The update process can be expressed as:
@@
    The $\epsilon$ represents a random sample drawn from $N(0,1)$ and MLP()($\cdot$) refers to a Multi-Layer Perceptron capable of doubling the dimension of $^{o}_{\hat{i}}$.
@@
VII WORLD: Tightening the Variational Lower Bound of Mutual Information Through Deconstructing the Problem Space
@@
The Brando network essentially functions as a distribution transformation mechanism, whose task is to map samples $o_{\hat{i}}$ drawn from a Gaussian distribution $N(\mu,\sigma^{2})$ into features $z_{\hat{i}j}$ of incorrect options customized for a given context matrix. This paper argues that instead of implicitly estimating the distribution of incorrect option features for each given context matrix, it is more effective to explicitly estimate the distribution of all RPM image features $z_{ij}$. In this way, when confronted with a specific context matrix, the required incorrect option features can be sampled in a targeted manner from this estimated distribution.
@@
VII-B Model the Distribution of RPM Image Features as a Gaussian Mixture Model.
@@
In order to enable the features $z_{ij}$ to be more easily projected onto the constructed GMM $\{U_{m}|m\in[1,M]\}$, the way DIO encodes features for RPM images requires minor modifications. The modifications are shown in Figure 7.
@@
Specifically, doubling the overall dimension of the ViT within DIO’s image feature extraction module can achieve the desired effect, causing the dimension of the latent features encoded for RPM images to double accordingly. After doubling the dimension, the resulting latent features are evenly split into two parts, denoted as $u_{ij}$ and $b_{ij}$ respectively. Then, the sum of $u_{ij}$ and $b_{ij}$ is taken as the new feature representation $z_{ij}$ for the image $x_{i}$. This process can be expressed by the following formula:
$\{u_{ij},b_{ij}\}=\text{ViT}(x_{i}),\,z_{ij}=u_{ij}+b_{ij}$ (34)
The new derivation method for $z_{ij}$ ensures that, when $b_{ij}$ follows a standard Gaussian distribution $\mathcal{N}(0,1)$, the distribution of $z_{ij}$ explicitly adheres to a Gaussian distribution $\mathcal{N}(u_{ij},1)$. Given this, we merely need to assign the latent features $u_{ij}$ of all observed RPM images onto specific components of the constructed GMM $\{U_{m}|m\in[1,M]\}$ and make the latent features $b_{ij}$ follow the distribution $N(0,1)$, so as to effectively model the distribution of $z_{ij}$ as a GMM with trainable means and known variances.
@@
    $\ell_{2}=\sum_{i=1}^{16}\sum_{j=1}^{N}\|b_{ij}-\epsilon_{ij}]\|_{2}^{2}$ (36)
    $\text{where}\quad\epsilon_{ij}\sim N(0,1)$
@@
$\ell_{\text{WORLD}}=\ell_{1}+\ell_{2}+\lambda\cdot\ell_{3}+\ell_{4}+\ell_{5}$ (40)
The coefficient $\lambda$ for $\ell_{3}$ is set to 0.25 in order to balance its relative magnitude compared to $\ell_{1}$.
By employing $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$ as the loss functions, we can train a DIO model where the distribution of image features $z_{ij}$ is modeled as a GMM. Notably, when training the DIO using $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, the feedforward mechanism undergoes the transformation illustrated in Figure 8.
@@
In Figure 8, the latent variable fed into the progressive pattern induction module within DIO is no longer $z_{ij}$, but rather $u_{ij}+\epsilon$. In other words, the feedforward process indicated by Equation (IV-C) is updated as follows:
$\{P_{jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}_{i=\alpha}$
@@
VII-C Leveraging the Modeled GMM to Provide Incorrect Options for Context Matrices
To steer DIO’s learning objective towards a tighter variational lower bound of the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, we also need to sample sufficient and diverse incorrect option features from the modeled GMM $\{U_{m}|m\in[1,M]\}$ for each given context matrix. Thanks to the inherent properties of GMM, targeted sampling within the set $\{U_{m}|m\in[1,M]\}$ is relatively straightforward. For a given context matrix $\{x_{i}|i\in[1,8]\}$, after identifying the components in the set $\{U_{m}|m\in[1,M]\}$ that correspond to the provided correct option $x_{\alpha}$, samples randomly drawn from the remaining components can be considered, to a certain extent, as features of incorrect options.
Identifying the components corresponding to the correct option, targeted sampling of incorrect option features, and steering the learning objective of DIO can all be achieved by introducing an additional loss function term:
${\ell_{\text{Weighted}}}=$ $-\sum_{j=1}^{N}\log\frac{{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{\tilde{m}}+\epsilon_{\tilde{m}}|i\in[1,8]\})}{|\tau|}}}}}{\sum_{m=1,m\neq\tilde{m}}^{M}{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{m}+\epsilon_{m}|i\in[1,8]\})}{|\tau|}}}}}$ (42)
where $\tilde{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{\alpha j}-{U}_{m}\|_{2}^{2},\,\epsilon_{m}\sim N(0,1)$
@@ VII-C Leveraging the Modeled GMM to Provide Incorrect Options for Context Matrices
The role of $\ell_{\text{Weighted}}$ is to fine-tune the DIO model. Specifically, when DIO approaches optimal performance under the constraint of the loss function $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, incorporating the additional loss term $\ell_{\text{Weighted}}$ to update the loss function to $\ell_{\text{WORLD}}+\ell_{\text{DIO}}+\ell_{\text{Weighted}}$ creates additional learning capacity and performance improvement potential. This adjustment enables the introduction of supplementary incorrect options during DIO’s late-stage training, thereby steering its learning objective toward a tighter variational lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, ultimately leading to further performance enhancement.

### v7

@@ V-B Tightening the Lower Bound via Hypothetical Options
Correspondingly, the variational lower bound on mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ is updated as follows:
$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq-\mathbb{E}\left[\log\frac{\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}{(\beta-9)\cdot e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}\right]$ (15)
Intuitively, this new lower bound is tighter. After incorporating the safeguard $\delta$, adopting it as the loss function for DIO should lead to improved performance.
However, this remains merely superficial. As shown in Equation (15), if these introduced hypothetical options fail to influence DIO’s decisions, this new lower bound, when used as an optimization objective, will not bring DIO any significant improvement. Specifically, when the probabilities that DIO assigns to the hypothetical options $\{x_{i}|i\in[17,\beta]\}$ approach zero, whereby these options cease to perturb the original distribution $\hat{P}(x_{{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)$, the variational lower bounds derived from Equation (11) and (15) become equivalent when employed as loss functions. This demonstrates that unconstructive options, which are readily assigned low weights and low probabilities by DIO, contribute nothing to tightening the lower bound.
In a nutshell, this paper proposes introducing extra constructive hypothetical options into DIO, such options that can reinforce acceptance of the correct options and strengthen repulsion of the incorrect ones, so as to steer its learning objective toward a tighter variational lower bound on mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ and, in turn, ultimately further enhance its performance.
VI Brando: Tightening the Variational Lower Bound on Mutual Information by Constructing Hypothetical Options.
This paper proposes incorporating additional constructive hypothetical options into DIO’s training process as a promising avenue for enhancing its performance. To this end, this section introduces a training framework, Brando (Bounded Random Options Mapping Method).
@@
The RPM problem expects the solving model to be trained with only limited supervision, such as the index of the correct option. Relying on manual design to create constructive options contradicts this intent; we therefore propose learning a set of constructive options instead.
We establish a set of trainable hypothetical options $\{x_{i}|i\in[17,\beta]\}$ for the context matrix $\{x_{i}|i\in[1,8]\}$. This paper is not concerned with the specific attributes of these options; rather, it focuses solely on how much these options can alter DIO’s evaluation of the original options $\{x_{i}|i\in[9,16]\}$. In other words, we are solely concerned with the constructiveness of these hypothetical options with respect to the context matrix $\{x_{i}|i\in[1,8]\}$ in DIO, rather than their specific configurations. To this end, we design a new loss function $\ell_{\text{Brando}}$ to train these hypothetical options for constructiveness:
$\ell_{\text{Brando}}=\ell_{\text{up}}+\frac{\ell_{down}}{\beta-17}$ (16)
where the loss term $\ell_{\text{up}}$:
$\ell_{\text{up}}=\log(\frac{\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}+\delta)-\log\delta$ (17)
and the loss term $\ell_{\text{down}}$:
$\ell_{\text{down}}=\sum_{{\hat{\beta}}=17}^{\beta}\log(\frac{\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}{e^{\text{DIO}(\{x_{i},x_{\hat{\beta}}|i\in[1,8]\})}}+\delta)-\log\delta$ (18)
@@
    The loss term $\ell_{\text{up}}$ requires DIO to be able to identify the correct option from a pool that combines the original candidate options $\{x_{i}|i\in[9,16]\}$ with the hypothetical options $\{x_{i}|i\in[17,\beta]\}$. It serves the purpose of tightening the variational lower bound on mutual information.
@@
    The loss term $\ell_{\text{down}}$ encourages hypothetical options $\{x_{i}|i\in[17,\beta]\}$ to attain higher probabilities than the original incorrect options $\{x_{i}|i\in[9,16],i\neq\alpha\}$ under DIO. It acts as a safeguard to prevent hypothetical options from degenerating into unconstructive ones.
By adopting $\ell_{\text{Brando}}$ as the training loss function, DIO not only obtains constructive hypothetical options but also parameterizes a tighter variational lower bound on $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
@@ VI-B Map Out Hypothetical Options
Clearly, the mapping $\mathcal{F}(\cdot)$ needs to yield structurally diverse and numerically flexible ($\beta-17$) hypothetical options conditioned on the context matrix $\{x_{i}|i\in[1,8]\}$, implying an inherent many-to-many property. This indicates that conventional deep neural networks struggle to fulfill the role of mapping $\mathcal{F}(\cdot)$. Firstly, constructing a network with variable and dynamic output specifications is highly challenging. Moreover, from the perspective of neural network optimization, achieving many-to-many mappings in a naive end-to-end fashion is infeasible. Given these challenges, we propose a many-to-one mapping as a substitute:
$x_{\hat{i}}=\hat{\mathcal{F}}(\{x_{i},o_{\hat{i}}|i\in[1,8]\}),\,o_{\hat{i}}\sim\mathcal{N}(\mu,\sigma^{2})$ (20)
@@
$z_{\hat{i}j}=\text{Brando}(\{z_{ij},o_{\hat{i}}|i\in[1,8]\}),\,o_{\hat{i}}\sim\mathcal{N}(\mu,\sigma^{2})$ (22)
Formula (22) shows that we abandon the option-mapping paradigm in favor of a representation-mapping approach, in order to reduce the computational load on Brando. The structure of the Brando network is shown in the Figure 5.
@@
As indicated in Figure 5, the Brando network is formed by stacking multiple logical layers, with no parameter sharing among each layer. The input to each logical layer consists of $\{z_{ij}|i\in[1,8]\}$ and $o_{\hat{i}}$. Except for the first layer, the input to all subsequent layers comes from the output of the preceding layer. For the first layer, $\{z_{ij}|i\in[1,8]\}$ are the features extracted by DIO from the context matrix, while the $o_{\hat{i}}$ is obtained through the reparameterization trick [2] from a learnable Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ parameterized by mean $\mu$ and variance $\sigma^{2}$. For the final layer, this paper treats its output $o_{\hat{i}}$ as the representation $z_{\hat{i}j}$ of a hypothetical option $x_{\hat{i}}$ provided by Brando with respect to the context matrix $\{x_{i}|i\in[1,8]\}$.
@@
    After these $Q$, $K$, and $V$ representations are appended with positional encoding and processed by layer normalization, the multi-head self-attention mechanism computes interactions among them, yielding the attention outputs $\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$:
@@
    As indicated by computation step 1 in Figure 5, the normalized $Q$ representations employ a linear attention mechanism to learn the historical input feature information stored in the memory bank $M_{n-1}$ and buffer $B_{n-1}$, resulting in the memory outputs $\{\tilde{z}_{ij},\tilde{o}_{\hat{i}}|i\in[1,8]\}$. The learning process can be interpreted as:
@@
    where the basis function $\phi(\cdot)$ is an element-wise ELU+1 operation.
@@
    As indicated by computation steps 2 and 3 in Figure 5, the memory bank $M_{n-1}$ and buffer $B_{n-1}$ are updated by leveraging the $K$ and $V$ representations derived from the current input features, an update process inspired by infinite attention [31] and expressed as:
@@
    The $\epsilon$ represents a random sample drawn from $\mathcal{N}(0,1)$ and MLP()($\cdot$) refers to a Multi-Layer Perceptron capable of doubling the dimension of $^{o}_{\hat{i}}$.
@@
VII WORLD: Tightening the Variational Lower Bound on Mutual Information Through Deconstructing the Problem Space
@@
The Brando network essentially functions as a distribution transformation mechanism, whose task is to map samples $o_{\hat{i}}$ drawn from a Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ into features $z_{\hat{i}j}$ of incorrect options customized for a given context matrix. This paper argues that instead of implicitly estimating the distribution of incorrect option features for each given context matrix, it is more effective to explicitly estimate the distribution of all RPM image features $z_{ij}$. In this way, when confronted with a specific context matrix, the required incorrect option features can be sampled in a targeted manner from this estimated distribution.
@@
VII-B Model the Distribution of RPM Image Features as a GMM.
@@
In order to enable the features $z_{ij}$ to be projected more easily onto the constructed GMM $\{U_{m}|m\in[1,M]\}$, the way DIO encodes features for RPM images requires minor modifications. The modifications are shown in Figure 7.
@@
Specifically, doubling the overall width of the ViT within DIO’s image feature extraction module can achieve the desired effect, causing the dimensions of the latent features encoded for RPM images to double accordingly. These doubled features are evenly split into two parts, denoted as $u_{ij}$ and $b_{ij}$. Then, their sum is taken as the new feature representation $z_{ij}$ for the image $x_{i}$. This updated module can be expressed by the following formula:
$\{u_{ij},b_{ij}\}=\text{WORLD-FEM}(x_{i}),\,z_{ij}=u_{ij}+b_{ij}$ (33)
The new derivation method for $z_{ij}$ ensures that, when $b_{ij}$ follows a standard Gaussian distribution $\mathcal{N}(0,1)$, the distribution of $z_{ij}$ explicitly adheres to a Gaussian distribution $\mathcal{N}(u_{ij},1)$. Given this, we merely need to assign the latent features $u_{ij}$ of all observed RPM images onto specific components of the constructed GMM $\{U_{m}|m\in[1,M]\}$ and make the latent features $b_{ij}$ follow the distribution $\mathcal{N}(0,1)$, so as to effectively model the distribution of $z_{ij}$ as a GMM with trainable means and known variances.
@@
    $\ell_{2}=\frac{1}{16\cdot N}\sum_{i=1}^{16}\sum_{j=1}^{N}\|b_{ij}-\epsilon_{ij}]\|_{2}^{2}$ (35)
    $\text{where}\quad\epsilon_{ij}\sim\mathcal{N}(0,1)$
@@
$\ell_{\text{WORLD}}=\ell_{1}+\ell_{2}+\ell_{3}+\lambda_{0}\cdot\ell_{4}+\lambda_{1}\cdot\ell_{5}$ (39)
Here, $\lambda_{0}$ and $\lambda_{1}$ are set to $5$ to balance the inductive difficulties between these terms.
By employing $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$ as the loss function, we can train a DIO model in which the distribution of image features $z_{ij}$ is modeled as a GMM. Notably, when training the DIO using $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, the feedforward mechanism undergoes the transformation illustrated in Figure 8.
@@
As shown in Figure 8, the latent variable fed into the progressive pattern induction module within DIO is no longer $z_{ij}$, but rather $u_{ij}+\epsilon_{ij}$. In other words, the feedforward process indicated by Equation (IV-C) is updated as follows:
$\{P_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$
@@
VII-C Targeted Sampling From the Learned GMM.
To steer DIO’s learning objective towards a tighter variational lower bound of the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, we also need to sample sufficient and diverse incorrect option features from the modeled GMM $\{U_{m}|m\in[1,M]\}$ for each given context matrix. Thanks to the inherent properties of GMM, targeted sampling within the set $\{U_{m}|m\in[1,M]\}$ is relatively straightforward. For a given context matrix $\{x_{i}|i\in[1,8]\}$, after identifying the components in the set $\{U_{m}|m\in[1,M]\}$ that correspond to the provided correct option $x_{\alpha}$, samples randomly drawn from the remaining components can be considered, to a certain extent, as features of the incorrect options.
Identifying the components corresponding to the correct option, targeted sampling of the incorrect option features, and steering the learning objective of DIO can all be achieved by introducing an additional loss function term:
${\ell_{\text{Weighted}}}$
$=\sum_{j=1}^{N}\log(\frac{\sum_{\begin{subarray}{c}m=1\\ m\neq\tilde{m}\end{subarray}}^{M}{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{m}+\epsilon_{m}|i\in[1,8]\})}{|\tau|}}}}}{{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{\tilde{m}}+\epsilon_{\tilde{m}}|i\in[1,8]\})}{|\tau|}}}}}+\delta)+\log\delta$ (41)
$\text{where}\quad\tilde{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{\alpha j}-{U}_{m}\|_{2}^{2},\,\,\epsilon_{m}\sim\mathcal{N}(0,1)$
@@
The role of $\ell_{\text{Weighted}}$ is to fine-tune the DIO model. Specifically, when DIO approaches optimal performance under the constraint of the loss function $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, incorporating the additional loss term $\ell_{\text{Weighted}}$ to update the loss function to $\ell_{\text{WORLD}}+\ell_{\text{DIO}}+\ell_{\text{Weighted}}$ creates additional learning capacity and potential of performance improvement. This adjustment enables the introduction of supplementary incorrect options during DIO’s late-stage training, thereby steering its learning objective toward a tighter variational lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, ultimately leading to further performance enhancement.